The Department of Computer Science and Engineering

University of Nevada, Reno

cordially invites you to a Master’s colloquium

CHIMP: The C/C++ Hybrid Imperative Meta-Programmer

A thesis submitted in partial fulfillment of the requirements for the degree of Master of Science with a major in Computer Science.

by

John L. Kenyon

Abstract: The C and C++ languages have been unchanged in over two decades, and during this time many shortcomings of the languages have become clear. Specifically, neither language supports reflection, and the meta-programming capabilities are very limited. Both of these problems can be solved by adding a pre-processing step, which can analyze and modify the code before the actual compiler translates it into object code. This means we can use metaprogramming and reflection to simplify some C/C++ tasks without having to change the languages at all.

2:00 pm, Thursday, March 13, 2008

Scrugham Engineering and Mines (SEM) room 201

For more information contact Dr. Fred Harris @ 784-6571 (Fred.Harris@cse.unr.edu)